# Hyper-SAMARL: Hypergraph-based Coordinated Task Allocation and Socially-aware Navigation for Multi-Robot Systems

### v1

@@ I Introduction
Multi-robot systems are becoming increasingly relevant in a wide range of real-world applications, including cleaning tasks in public spaces [1, 2, 3, 4, 5] like airports, monitoring operations in large facilities, and delivery robots on university campuses. In these scenarios, a team of robots must navigate shared spaces that are often filled with humans, such as pedestrians, in order to reach assigned points of interest (POIs) and complete specific tasks. As these systems become more prevalent, ensuring that robots can safely and efficiently navigate these human-filled environments while maintaining socially acceptable behavior presents a critical challenge.
@@
Figure 1: An illustration of multi-robot task allocation and social navigation: mobile robots are engaging in cooperative navigation toward POIs in a dynamic, human-filled environment while adhering to social norms.
This study addresses the problem of multi-robot navigation in complex, human-populated environments, as illustrated in Fig. 1. The core challenge lies in autonomously assigning POIs to multiple robots while determining feasible and socially-aware paths that allow the robots to navigate around humans. Although task allocation for multi-robot systems has been well explored [17, 18, 19, 20, 21, 22, 23], many existing approaches assume static obstacle configurations and struggle to adapt to the dynamic and unpredictable nature of human environments. Human movement frequently alters the landscape in real time, making static strategies insufficient for allowing robots to effectively reach their destinations. To overcome these limitations, we extend socially-aware navigation algorithms [6] to improve human-robot cooperation, with a specific focus on enabling multi-robot systems to operate effectively in dynamic, human-filled public spaces.
@@ II Related Works
More recent work focuses on learning-based approaches that couple HRI inference with path planning [7, 6, 33, 34], leveraging neural networks or preference distribution to approximate human preferences and cooperation tendencies. For multi-agent scenarios, studies such as [35, 36] introduced training paradigms that integrate multi-robot systems into a single policy network using RL algorithms like PPO. More recently, [37, 6] extended single-agent RL to multi-agent reinforcement learning (MARL) with the MARL benchmark multi-agent proximal policy optimization (MAPPO) algorithm to train graph-based or transformer-based neural networks for multi-robot social navigation. Despite these promising developments in implicit coordination for multi-robot social navigation, the integration of advanced task allocation and navigation strategies remains underexplored. To address this gap, we frame the multi-robot socially-aware navigation (MR-TASN) task as a decentralized partially observable semi-Markov decision process (Dec-POSMDP) problem [38] to couple task allocation with social navigation.
Path planning and task allocation in multi-robot systems are inherently complex, requiring coordination and self-learning in dynamic and uncertain environments. Several approaches have been proposed to tackle these challenges. For instance, [39] introduced a cluster-based task allocation approach method grounded in game theory. [40] formulated the multi-robot task allocation problem as an optimization task using a graph structure. Also, [41] constructed a hypergraph to represent the correlations between robots and objects in the environment. However, many of these approaches assume statistic obstacle configurations [42] or fail to address the adaptability of task re-assignment in dynamic, human-filled environments.
In this work, we explore the feasibility of multi-robot task allocation in dynamic, human-populated environments, framing it as an MR-TASN problem. We leverage hypergraph diffusion to dynamically adjust task allocations in response to real-time environmental changes. Our proposed framework, Hyper-SAMARL, provides a more adaptive and robust solution to the challenges of task allocation and socially-aware navigation in multi-robot systems.
@@ III-B Markov Decision Process Formulation
Drawing inspiration from [6, 38], the MR-TASN task is formally constructed as a Dec-POSMDP with the tuple $\langle\mathcal{S},\mathcal{U},\mathcal{A},\mathcal{O},\mathcal{P},\mathcal{R},\mathbf{R},\mathcal{C},\mathcal{S}_{0},\gamma\rangle$. The joint state is $\hat{s}_{t}=[s_{t}^{r_{1}},\cdots,s_{t}^{r_{n}}]\in\mathcal{S}$, and the joint observation is $\hat{o}_{t}=[o_{t}^{r_{1}},\cdots,o_{t}^{r_{n}}]\in\mathcal{O}$. Hyper-SAMARL converts local observations into macro-action (MA) $\hat{u}=[p_{gx}^{(\cdot)},p_{gy}^{(\cdot)}]\in\mathcal{U}$ via robot policy $\hat{\pi}$ to assign POI positions $\{p_{gx},p_{gy}\}$ and guide the generation of velocity-based local-action (LA) $\hat{a}=[v_{x}^{(\cdot)},v_{y}^{(\cdot)}]\in\mathcal{A}$. Specifically, the robots’ MA are updated at each decision-making timestep $t_{k}\in\{t_{0},\cdots,t_{K}\}$, while the robots’ LA are sequentially executed during the intervals between decision-making timesteps $t\in[t_{k},t_{k+1})$.
@@ IV-B Hypergraph Diffusion
The hypergraph of MR-TASN is initialized to explicitly represent the correlation between multiple objectives, based on spatial Euclidean distance similarity [43]. Formally, the MR-TASN hypergraph is initialized by grouping spatially closest objectives (robots, pedestrians, and POIs), which are connected as hyperedges. In the initialized hypergraph, the environmental dynamics of the social context are encoded into hypergraph vertex features, while the spatial distance information between objectives is stored in the hyperedges. Once the hypergraph of MR-TASN is constructed, we embed the vertex features based on the spatial-temporal dependencies. Due to the lack of exhaustive HRI relational reasoning in the hypergraph initialization, Hyper-SAMARL employs a hypergraph diffusion network to propagate and reorganize the vertex features and hypergraph structure, facilitating effective task allocation and smooth, socially acceptable behavior execution, as illustrated in Fig. 2.
@@ 19                         Update macro action ${{u}_{t_{k}}^{i}}\sim\pi_{\theta}^{i}(\mathcal{G}_{t_{k}}^{i,*}$);
22                  Execute $\mathrm{a}_{t}^{i}\sim{\pi}_{\theta^{\prime}}^{i}(o_{t_{k}}^{i},{{u}_{t_{k}}^{i}};\theta^{\prime},\mathbf{X}_{ST},H_{t_{k}}^{i})$;
@@ 29      Update $(\theta,\theta^{\prime},\psi,\psi^{\prime},p_{star},{p}_{diff})$ with MAPPO;
where $f_{t}(\cdot)=0.15\cdot t$ represents the penalty for algorithm time efficiency.
@@ V-A Simulation Environment
To address the MR-TASN task, we designed several simulated scenarios where a group of robots is assigned to search for individual target POIs in a human-filled environment. Formally, the MR-TASN task is modeled as a Dec-POSMDP, which couples the execution of temporally-extended macro actions for task allocation with the generation of primitive actions (LAs) for social navigation. The objective $\mathcal{J}$ and the blended MA&LA policies are optimized using a MARL method, utilizing a hypergraph-based neural network.
@@ V-A4 Results
Notably, our proposed framework, Hyper-SAMARL, outperforms other baselines and ablation models in both the AS and SS metrics, as well as in reward collection, as shown in Table I and Fig. 3. The use of a hypergraph-based neural network not only facilities the inference and understanding of complex environmental dynamics and potential correlations between objects but also enhances the adaptability of task assignments across different environmental conditions. Additionally, the cooperative strategies of the multi-robot system are integrated into the hypergraph-based neural network, developed through the MARL CTDE training procedure. In summary, the experimental results demonstrate that Hyper-SAMARL provides a benchmark-level performance for addressing the MR-TASN task in dynamic environments. Videos of the experiments on the MR-TASN task can be found in the supplementary material.

### v2

# Hypergraph-based Coordinated Task Allocation and Socially-aware Navigation for Multi-Robot Systems

@@ I Introduction
Multi-robot systems are becoming increasingly relevant in a wide range of real-world applications, including cleaning tasks in public spaces [1, 2, 3, 4, 5] like airports, monitoring operations in large facilities, and delivery robots on university campuses. In these scenarios, a team of robots must navigate shared spaces that are often filled with humans, such as pedestrians, in order to reach assigned points of interest (POIs) and complete specific tasks. As these systems become more prevalent, ensuring that robots can safely and efficiently navigate these human-filled environments while maintaining socially acceptable behavior [6, 7, 8] presents a critical challenge.
@@
Figure 1: An illustration of multi-robot task allocation and social navigation task and simulator: mobile robots are engaging in cooperative navigation toward POIs in a human-filled environment while adhering to social norms.
This study addresses the problem of multi-robot navigation in complex, human-populated environments, as illustrated in Fig. 1. The core challenge lies in autonomously assigning POIs to multiple robots while determining feasible and socially-aware paths that allow the robots to navigate around humans. Although task allocation for multi-robot systems has been well explored [20, 21, 22, 23, 24, 25, 26], many existing approaches assume static obstacle configurations [27, 28, 29] and struggle to adapt to the dynamic and unpredictable nature of human environments. Human movement frequently alters the landscape in real time, making static strategies insufficient for allowing robots to effectively reach their destinations. To overcome these limitations, we extend socially-aware navigation algorithms [9] to improve human-robot cooperation, with a specific focus on enabling multi-robot systems to operate effectively in dynamic, human-filled public spaces.
@@ II Related Works
More recent work focuses on learning-based approaches that couple HRI inference with path planning [10, 9, 38], leveraging neural networks or preference distribution to approximate human preferences and cooperation tendencies. For multi-agent scenarios, studies such as [39, 40] introduced training paradigms that integrate multi-robot systems into a single policy network using RL algorithms like PPO. More recently, [41, 9] extended single-agent RL to multi-agent reinforcement learning (MARL) with the MARL benchmark multi-agent proximal policy optimization (MAPPO) algorithm to train graph-based or transformer-based neural networks for multi-robot social navigation. Despite these promising developments in implicit coordination for multi-robot social navigation, the integration of advanced task allocation and navigation strategies remains underexplored. To address this gap, we frame the multi-robot socially-aware navigation (MR-TASN) task as a decentralized partially observable semi-Markov decision process (Dec-POSMDP) problem [42] to couple task allocation with social navigation.
Path planning and task allocation in multi-robot systems are inherently complex, requiring coordination and self-learning in dynamic and uncertain environments. Several approaches have been proposed to tackle these challenges. For instance, [43] introduced a cluster-based task allocation approach method grounded in game theory. [44] formulated the multi-robot task allocation problem as an optimization task using a graph structure. Also, [45] constructed a hypergraph to represent the correlations between robots and objects in the environment. However, many of these approaches assume static obstacle configurations [27, 28, 29] or fail to address the adaptability of task re-assignment in dynamic, human-filled environments.
In this work, we explore the feasibility of multi-robot task allocation in dynamic, human-populated environments, framing it as an MR-TASN problem. We leverage hypergraph diffusion to dynamically adjust both task allocations and navigation behaviors in response to real-time environmental changes. Our proposed framework, Hyper-SAMARL, provides a more adaptive and robust solution to the challenges of task allocation and socially-aware navigation in multi-robot systems.
@@ III-B Markov Decision Process Formulation
Drawing inspiration from [9, 42], the MR-TASN task is formally constructed as a Dec-POSMDP with the tuple $\langle\mathcal{S},\mathcal{U},\mathcal{A},\mathcal{O},\mathcal{P},\mathcal{R},\mathbf{R},\mathcal{C},\mathcal{S}_{0},\gamma\rangle$. The joint state is $\hat{s}_{t}=[s_{t}^{r_{1}},\cdots,s_{t}^{r_{n}}]\in\mathcal{S}$, and the joint observation is $\hat{o}_{t}=[o_{t}^{r_{1}},\cdots,o_{t}^{r_{n}}]\in\mathcal{O}$. Hyper-SAMARL converts local observations into macro-action (MA) $\hat{u}=[p_{gx}^{(1,\cdots,n)},p_{gy}^{(1,\cdots,n)}]\in\mathcal{U}$ via robot policy $\hat{\pi}$ to assign POI positions $\{p_{gx},p_{gy}\}$ and guide the generation of velocity-based local-action (LA) $\hat{a}=[v_{x}^{(1,\cdots,n)},v_{y}^{(1,\cdots,n)}]\in\mathcal{A}$. Specifically, the robots’ MA are updated at each decision-making timestep $t_{k}\in\{t_{0},\cdots,t_{K}\}$, while the robots’ LA are sequentially executed during the intervals between decision-making timesteps $t\in[t_{k},t_{k+1})$.
@@ IV-B Hypergraph Diffusion
The hypergraph of MR-TASN is initialized to explicitly represent the correlation between multiple objectives, based on spatial Euclidean distance similarity [46]. Formally, the MR-TASN hypergraph is initialized by grouping spatially closest objectives (robots, pedestrians, and POIs), which are connected as hyperedges. In the initialized hypergraph, the environmental dynamics of the social context are encoded into hypergraph vertex features, while the spatial distance information between objectives is stored in the hyperedges. Once the hypergraph of MR-TASN is constructed, we embed the vertex features based on the spatial-temporal dependencies $\mathbf{X}_{ST}$. Due to the lack of exhaustive HRI relational reasoning in the hypergraph initialization, Hyper-SAMARL employs a hypergraph diffusion network to propagate and reorganize the vertex features and hypergraph structure, facilitating effective task allocation and smooth, socially acceptable behavior execution, as illustrated in Fig. 2.
@@ 19                         Update macro action ${{u}_{t_{k}}^{i}}\sim\pi_{\theta}^{i}(\mathcal{G}_{t_{k}}^{i,*}$);
22                  Execute $\mathrm{a}_{t}^{i}\sim{\pi}_{\theta^{\prime}}^{i}(o_{t_{k}}^{i},{{u}_{t_{k}}^{i}};\theta^{\prime},\mathcal{G}_{t_{k}}^{i,*},H_{t_{k}}^{i})$;
@@ 29      Update $(\theta,\theta^{\prime},\psi,\psi^{\prime},p_{star},{p}_{diff})$ with MAPPO;
where $f_{t}(\cdot)=\kappa_{t}\cdot t$ represents the penalty for algorithm time efficiency with respect to the penalty factor $\kappa_{t}\in[0,1]$.
@@ V-A Simulation Environment
To address the MR-TASN task, we designed several simulated scenarios where a group of robots is assigned to search for individual target POIs in a human-filled environment, as shown in Fig. 1. Formally, the MR-TASN task is modeled as a Dec-POSMDP, which couples the execution of temporally-extended macro actions for task allocation with the generation of primitive actions (LAs) for social navigation. The objective $\mathcal{J}$ and the blended MA&LA policies are optimized using a MARL method, using a hypergraph-based network.
@@ V-A4 Results
Notably, our proposed framework, Hyper-SAMARL, outperforms other baselines and ablation models in both the AS and SS metrics, as well as in reward collection, as shown in Table I and Fig. 3. The use of a hypergraph-based neural network not only facilities the inference and understanding of complex environmental dynamics and potential correlations between objects but also enhances the adaptability of task assignments across different environmental conditions. Additionally, the cooperative strategies of the multi-robot system are integrated into the hypergraph-based neural network, developed through the MARL CTDE training procedure. In summary, the experimental results demonstrate that Hyper-SAMARL provides a benchmark-level performance for addressing the MR-TASN task in dynamic environments. Videos of the experiments on the MR-TASN task can be found https://sites.google.com/view/hyper-samarl.